OpenStudy (anonymous):

Can you explain how to do this to refresh my memory? 4(x - 1) = 2(x + 2) + 6

OpenStudy (anonymous):

dont you just multiply the 4(x-1) and the 2(x+2) then combine like terms

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

4(x-1)=2(x+2)+6 4x-4=2x+4+6 4x-4=2x+10 +4 +4 4x=2x+14 -2x -2x 2x=14 divide everything by 2 to get the x to be alone. 14 divided by 2 is 7. therefore, x=7

hartnn (hartnn):

and to make things easier, you can also divide everything by 2 from the 1st step itself.

OpenStudy (anonymous):

yeah, thats right, thatll make it wayyy easier

OpenStudy (anonymous):

Okay thank you!